The sum of the speeds of two trains is 723.7. If the speed of the first train is 10.3 faster than the second train, find the speeds of each.
HELP!! this is killing me that i cant figure this out

Answer by Maths68(1474)

Let
Speed of slower train = x
Speed of faster train = x+10.3

Speed of slower train + Speed of faster train = Sum of speeds
x+x+10.3=723.7
2x+10.3=723.7
2x=723.7-10.3
2x=713.4
2x/2=713.4/2
x=356.7


Speed of slower train = x = 356.7
Speed of faster train = x+10.3 = 356.7+10.3=367